# Convexified Message-Passing Graph Neural Networks

### v1

@@ Appendix E More Details on Matrix Approximation and Factorization Methods
where we used the triangle inequality and the fact that the singular vectors are unit vectors. Since $[\Psi^{\mathcal{A}_{\ell}}(\mathcal{X})]_{if}=\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, we obtained that $\mathcal{F}_{\mathcal{B}_{\ell}}\subset\mathcal{F}_{\text{cgcn}}$, as desired.
To conclude the proof, we next prove that any empirical risk minimizer $\Phi^{\mathcal{A}}$ in $\mathcal{F}_{\text{cgcn}}$ is also in $\mathcal{F}_{\mathcal{B}_{\ell}}$. By our proof in Appendix D, the value of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ is then given by $\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}^{(j)}))$, where, as specified in (25), $\tau_{\ell}^{f}(\mathcal{Z})=\sum_{k=0}^{K}\langle(Q_{\ell}^{k})^{\dagger}\mathbf{v}_{\ell}^{k}(\mathbf{z}_{k}),(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\rangle$ is a reparameterization of the graph filters at layer $\ell$ for some vector of the coefficients $\mathbf{\beta}_{\ell k}^{f}\in\mathbb{R}^{nm}$ whose $(i,j)$-th coordinate is $\beta_{\ell k,(i,j)}^{f}$. That is, $(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ now act as the filters under this reparameterization. In Appendix D, we have shown that the Hilbert norm of $\tau_{\ell}^{f}$ is then $\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}=\|(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\|_{2}\leq C_{\sigma,\ell}(R_{\ell})$, where the last inequality is by (22). By (26), we further infer that the nuclear norm of the filters $(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ in the new parameterization satisfies $\|(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\|_{*}\leq C_{\sigma,\ell}(R_{\ell})\sqrt{F_{\ell}F_{\ell-1}(K+1)}\leq C_{\sigma,\ell}(R_{\ell})(K+1)\sqrt{F_{\ell}F_{\ell-1}}$, meaning that $\Phi^{\mathcal{A}}\in\mathcal{F}_{\mathcal{B}_{\ell}}$, as desired. ∎
Next, we prove that $\mathcal{F}_{L,cgcn}$ is not "too big" by upper bounding its Rademacher complexity. The Rademacher complexity is a key concept in empirical process theory and can be used to bound the generalization error in our empirical risk minimization problem. Readers should refer to [2] for a brief overview on Rademacher complexity. The Rademacher complexity of a function class =$\mathcal{F}=\{\phi:\mathcal{D}\rightarrow\mathbb{R}\}$ with respect to $m$ i.i.d samples $\{\mathcal{X}^{(j)}\}_{j=1}^{m}$ is given by:
@@
where $\{\varepsilon_{j}\}_{j=1}^{m}$ are an i.i.d. sequence of uniform {±$\{\pm 1\}$-valued variables. Next, we provide the upper bound on the Rademacher complexity of the function class $\mathcal{F}_{L,cgcn}$ describing the graph filters. The lemma refers to the random kernel matrix $\mathcal{K}_{\ell}^{k}(\mathcal{X})\in\mathbb{R}^{n\times n}$ induced by a graph signal $\mathcal{X}\in\mathbb{R}^{n\times F}$ drawn randomly from the population, i.e., the $(i,i^{\prime})$-th entry of $\mathcal{K}_{\ell}^{k}(\mathcal{X})$ is $\kappa_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}),\mathbf{z}_{i^{\prime}}^{k}(\mathcal{X}))$. Furthermore, we consider the expectation $\mathbb{E}[||\mathcal{K}_{\ell}^{k}(\mathcal{X})||_{2}]$ of this matrix’s spectral norm. While the previous lemmas consider the most general case where the number of features $F_{L}=G$ at the last layer may be any positive integer, observe that the following result focuses on the the binary classification case where $F_{L}=G=1$. As mentioned in Section 4.5, Lemma F.3 can be readily extended to the multi-class case by employing a standard one-versus-all reduction to the binary case.
Lemma F.3.
In the binary classification case where $F_{L}=G=1$, there exists a universal constant $c>0$ such that:
$\mathcal{R}_{m}(\mathcal{F}_{L,cgcn})\leq\frac{c\cdot C_{\sigma,\ell}(\mathcal{B}_{L})\sqrt{F_{\ell-1}(K+1)\cdot\log(n(K+1))\cdot\sum_{k=0}^{K}\mathbb{E}[||K_{L}^{k}(\mathcal{X})||_{2}]}}{\sqrt{m}}$ (36)
@@
We begin with some preliminaries that apply to the general case and then leverage them in order to obtain the result. For brevity, we denote $R_{\ell}=C_{\sigma,\ell}(\mathcal{B}_{\ell})\sqrt{F_{\ell}F_{\ell-1}}$. Consider some function $\Phi^{\mathcal{A}}=\phi^{{\mathcal{A}}_{L}}\circ\dots\circ\phi^{{\mathcal{A}}_{1}}\in\mathcal{F}_{\text{cgcn}}$. Without loss of generality, we assume that, each input input graph signal $\mathcal{X}^{(j)}$, the feature vector $[\mathcal{X}^{(j)}]_{i}=\mathbf{x}_{i}^{(j)}$ of each node $i$ resides in the unit $\ell_{2}$-ball (which can be obtain via normalization). Combined with $\kappa_{\ell}$’s being a continuous kernel with $\kappa_{\ell}(\mathbf{z},\mathbf{z})\leq 1$, Mercer’s theorem [44, Theorem 4.49] implies that there exists a feature map $\varphi_{\ell}:\mathbb{R}^{F_{\ell-1}}\rightarrow\ell^{2}(\mathbb{N})$ for which $\sum_{\xi=1}^{\infty}[\varphi_{\ell}(\mathbf{z})]_{\xi}[\varphi_{\ell}(\mathbf{z}^{\prime})]_{\xi}=\sum_{\xi=1}^{\infty}\varphi_{\ell\xi}(\mathbf{z})\varphi_{\ell\xi}(\mathbf{z}^{\prime})$ converges uniformly and absolutely to $\kappa_{\ell}(\mathbf{z},\mathbf{z}^{\prime})$. Accordingly, it holds that $\kappa_{\ell}(\mathbf{z},\mathbf{z}^{\prime})=\langle\varphi_{\ell}(\mathbf{z}),\varphi_{\ell}(\mathbf{z}^{\prime})\rangle$, which yields the following for each feature $f$ of any node $i$ at layer $\ell$ is $\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, where for any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is given by $[\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for any $k\in[K]\cup\{0\}$:
@@
where $\bar{\mathbf{a}}^{f}_{\ell k}\in\ell^{2}(\mathbb{N})$ is a countable-dimensional vector, due to the fact that $\varphi$ itself is a countable sequence of functions. Particularly, the Hilbert norm of $\tau_{\ell}^{f}$ is $\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}=\|\bar{\mathbf{a}}^{f}_{\ell k}\|_{2}$. Now, for any graph signal $\mathcal{X}$, let $\Theta_{i}(\mathcal{X})$ be the linear operator that maps any sequence $\theta\in\ell^{2}(\mathbb{N})$ to the vector $[\langle\varphi_{\ell}(\mathbf{z}_{i}^{0}(\mathcal{X})),\theta\rangle,\dots,\langle\varphi_{\ell}(\mathbf{z}_{i}^{K}(\mathcal{X})),\theta\rangle]^{\top}\in\mathbb{R}^{K+1}$. Thereby, recalling that $\Phi^{\mathcal{A}}=\phi^{{\mathcal{A}}_{L}}\circ\dots\circ\phi^{{\mathcal{A}}_{1}}$ where $\phi^{{\mathcal{A}}_{\ell}}\in\mathcal{F}_{\ell,cgcn}$ satisfies $[\phi^{{\mathcal{A}}_{\ell}}(\mathcal{X}_{\ell-1})]_{if}=\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, then the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ can be rephrased as:
=${}_{if}=\mathbf{1}_{K+1}^{\top}\Theta_{i}(\mathcal{X}_{\ell-1})\bar{\mathbf{a}}^{f}_{\ell k}=\operatorname*{trace}\Big{(}\Theta_{i}(\mathcal{X}_{\ell-1})(\bar{\mathbf{a}}^{f}_{\ell k}\mathbf{1}_{K+1}^{\top})\Big{)}$ (38)
@@
Now, we obtain our upper bound for the binary classification case where $F_{L}=G=1$. Combining (38) with (39), we obtain that the Rademacher complexity of $\mathcal{F}_{L,cgcn}$ is upper bounded as follows:
@@
Since $\sum_{\xi=1}^{\infty}(\varphi_{\ell\xi}(\mathbf{z}))^{2}$ uniformly converges to $\kappa_{\ell}(\mathbf{z})$, the second term in (41c) converges to 00 as $r\rightarrow\infty$. Hence, it is sufficient to upper bound the first term in (41c) and take the limit $r\rightarrow\infty$. By Bernstein inequality due to [30, Theorem 2.1], whenever $\operatorname*{trace}(\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})^{\top})\leq C_{1}$, there is a universal constant $c>0$ such that the expected spectral norm is upper bounded by:
@@
We are ready to prove the key lemma that will yield our main result. Recall that our current focus is on binary classification (i.e., $F_{L}=G=1$). While our prior results consider models with multiple hidden layers, our main result regards the case of a single hidden layer $L=1$. As mentioned in Section 4.5, in the case of multiple hidden layers, our result applies to each individual layer separately. Using Lemmas F.1F.3, we are capable of comparing the CGCN predictor obtained by Algorithm 2 against optimal model in the GCN class. Lemma F.2 yields that the predictor $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\cdot)$ generated at the last layer is an empirical risk minimizer in the function class $\mathcal{F}_{\text{cgcn}}$. As such, by the theory of Rademacher complexity (see, e.g., [2]), we obtain that:
@@
where $c$ is a universal constant a and we used the assumption that $J$ is $M$-Lipschitz. By Lemma F.1, we have that $\inf_{\Phi\in\mathcal{F}_{L,cgcn}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]\leq\inf_{\Phi\in\mathcal{F}_{L,gcn}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]$. Since $\mathcal{F}_{\text{cgcn}}=\mathcal{F}_{1,cgcn}$ due to $L=1$, this means that $\inf_{\Phi\in\mathcal{F}_{\text{cgcn}}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]\leq\inf_{\Phi\in\mathcal{F}_{\text{gcn}}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]$. Combined with (43), we obtain the desired.
Remark F.4.
The constant $C_{\sigma,L}(\mathcal{B}_{L})$ depends on the convergence rate of the polynomial expansion of the activation function $\sigma$ (See Appendix B). Algorithmically, knowing the activation function is not required to run Algorithm 2. Theoretically, however, the choice of $\sigma$ matters for Theorem 4.7 when comparing the predictor produced by Algorithm 2 against the best GCN, as it affects the representation power of such GCN. If a GCN with an activation function $\sigma$ performs well, CGCN will similarly generalize strongly.
Appendix G Additional Experimental Details
G.1 Dataset Descriptions
Table 2 provides a summary of the statistics and characteristics of datasets used in this paper. All datasets are accessed using the TUDataset class [31] in PyTorch Geometric [14] (MIT License).

### v2

@@ Appendix F MORE DETAILS ON MATRIX APPROXIMATION AND FACTORIZATION METHODS
where we used the triangle inequality and the fact that the singular vectors are unit vectors. Since $\Psi^{\mathcal{A}_{\ell}}(\mathcal{X})]_{if}=\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, we obtained that $\mathcal{F}_{\mathcal{B}_{\ell}}\subset\mathcal{F}_{\text{cgcn}}$, as desired.
To conclude the proof, we next prove that any empirical risk minimizer $\Phi^{\mathcal{A}}$ in $\mathcal{F}_{\text{cgcn}}$ is also in $\mathcal{F}_{\mathcal{B}_{\ell}}$. By our proof in Appendix E, the value of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ is then given by $\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}^{(j)}))$, where, as specified in (25), $\tau_{\ell}^{f}(\mathcal{Z})=\sum_{k=0}^{K}\langle(Q_{\ell}^{k})^{\dagger}\mathbf{v}_{\ell}^{k}(\mathbf{z}_{k}),(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\rangle$ is a reparameterization of the graph filters at layer $\ell$ for some vector of the coefficients $\mathbf{\beta}_{\ell k}^{f}\in\mathbb{R}^{nm}$ whose $(i,j)$-th coordinate is $\beta_{\ell k,(i,j)}^{f}$. That is, $(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ now act as the filters under this reparameterization. In Appendix E, we have shown that the Hilbert norm of $\tau_{\ell}^{f}$ is then $\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}=\|(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\|_{2}\leq C_{\sigma,\ell}(R_{\ell})$, where the last inequality is by (22). By (26), we further infer that the nuclear norm of the filters $(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ in the new parameterization satisfies $\|(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\|_{*}\leq C_{\sigma,\ell}(R_{\ell})\sqrt{F_{\ell}F_{\ell-1}(K+1)}\leq C_{\sigma,\ell}(R_{\ell})(K+1)\sqrt{F_{\ell}F_{\ell-1}}$, meaning that $\Phi^{\mathcal{A}}\in\mathcal{F}_{\mathcal{B}_{\ell}}$, as desired. ∎
Next, we prove that $\mathcal{F}_{L,cgcn}$ is not ”too big” by upper bounding its Rademacher complexity. The Rademacher complexity is a key concept in empirical process theory and can be used to bound the generalization error in our empirical risk minimization problem. Readers should refer to (Bartlett and Mendelson, 2002) for a brief overview on Rademacher complexity. The Rademacher complexity of a function class $\mathcal{F}=\{\phi:\mathcal{D}\rightarrow\mathbb{R}\}$ with respect to $m$ i.i.d samples $\{\mathcal{X}^{(j)}\}_{j=1}^{m}$ is given by:
@@
[…]
Lemma G.3.
In the binary classification case where $F_{L}=G=2$, there exists a universal constant $c>0$ such that:
[…]
@@
We begin with some preliminaries that apply to the general case and then leverage them in order to obtain the result. For brevity, we denote $R_{\ell}=C_{\sigma,\ell}(\mathcal{B}_{\ell})\sqrt{F_{\ell}F_{\ell-1}}$. Consider some function $\Phi^{\mathcal{A}}=\phi^{{\mathcal{A}}_{L}}\circ\dots\circ\phi^{{\mathcal{A}}_{1}}\in\mathcal{F}_{\text{cgcn}}$. Without loss of generality, we assume that, each input input graph signal $\mathcal{X}^{(j)}$, the feature vector $\mathcal{X}^{(j)}]_{i}=\mathbf{x}_{i}^{(j)}$ of each node $i$ resides in the unit $\ell_{2}$-ball (which can be obtain via normalization). Combined with $\kappa_{\ell}$’s being a continuous kernel with $\kappa_{\ell}(\mathbf{z},\mathbf{z})\leq 1$, Mercer’s theorem (Steinwart, 2008, Theorem 4.49) implies that there exists a feature map $\varphi_{\ell}:\mathbb{R}^{F_{\ell-1}}\rightarrow\ell^{2}(\mathbb{N})$ for which $\sum_{\xi=1}^{\infty}[\varphi_{\ell}(\mathbf{z})]_{\xi}[\varphi_{\ell}(\mathbf{z}^{\prime})]_{\xi}=\sum_{\xi=1}^{\infty}\varphi_{\ell\xi}(\mathbf{z})\varphi_{\ell\xi}(\mathbf{z}^{\prime})$ converges uniformly and absolutely to $\kappa_{\ell}(\mathbf{z},\mathbf{z}^{\prime})$. Accordingly, it holds that $\kappa_{\ell}(\mathbf{z},\mathbf{z}^{\prime})=\langle\varphi_{\ell}(\mathbf{z}),\varphi_{\ell}(\mathbf{z}^{\prime})\rangle$, which yields the following for each feature $f$ of any node $i$ at layer $\ell$ is $\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, where for any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is given by $\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for any $k\in[K]\cup\{0\}$:
@@
where $\bar{\mathbf{a}}^{f}_{\ell k}\in\ell^{2}(\mathbb{N})$ is a countable-dimensional vector, due to the fact that $\varphi$ itself is a countable sequence of functions. Particularly, the Hilbert norm of $\tau_{\ell}^{f}$ is $\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}=\|\bar{\mathbf{a}}^{f}_{\ell k}\|_{2}$. Now, for any graph signal $\mathcal{X}$, let $\Theta_{i}(\mathcal{X})$ be the linear operator that maps any sequence $\theta\in\ell^{2}(\mathbb{N})$ to the vector $\langle\varphi_{\ell}(\mathbf{z}_{i}^{0}(\mathcal{X})),\theta\rangle,\dots,\langle\varphi_{\ell}(\mathbf{z}_{i}^{K}(\mathcal{X})),\theta\rangle]^{\top}\in\mathbb{R}^{K+1}$. Thereby, recalling that $\Phi^{\mathcal{A}}=\phi^{{\mathcal{A}}_{L}}\circ\dots\circ\phi^{{\mathcal{A}}_{1}}$ where $\phi^{{\mathcal{A}}_{\ell}}\in\mathcal{F}_{\ell,cgcn}$ satisfies $\phi^{{\mathcal{A}}_{\ell}}(\mathcal{X}_{\ell-1})]_{if}=\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, then the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ can be rephrased as:
${}_{if}=\mathbf{1}_{K+1}^{\top}\Theta_{i}(\mathcal{X}_{\ell-1})\bar{\mathbf{a}}^{f}_{\ell k}=\operatorname*{trace}\Big(\Theta_{i}(\mathcal{X}_{\ell-1})(\bar{\mathbf{a}}^{f}_{\ell k}\mathbf{1}_{K+1}^{\top})\Big)$ (38)
@@
Now, we obtain our upper bound for the binary classification case where $F_{L}=G=2$. Combining (38) with (39), we obtain that the Rademacher complexity of $\mathcal{F}_{L,cgcn}$ is upper bounded as follows:
@@
Since $\sum_{\xi=1}^{\infty}(\varphi_{\ell\xi}(\mathbf{z}))^{2}$ uniformly converges to $\kappa_{\ell}(\mathbf{z})$, the second term in (41c) converges to 0 as $r\rightarrow\infty$. Hence, it is sufficient to upper bound the first term in (41c) and take the limit $r\rightarrow\infty$. By Bernstein inequality due to (Minsker, 2017, Theorem 2.1), whenever $\operatorname*{trace}(\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})^{\top})\leq C_{1}$, there is a universal constant $c>0$ such that the expected spectral norm is upper bounded by:
@@
We are ready to prove the key lemma that will yield our main result. Recall that our current focus is on binary classification (i.e., $F_{L}=G=2$). While our prior results consider models with multiple hidden layers, our main result regards the case of a single hidden layer $L=1$. As mentioned in Section 4.5, in the case of multiple hidden layers, our result applies to each individual layer separately. Using Lemmas G.1G.3, we are capable of comparing the CGCN predictor obtained by Algorithm 2 against optimal model in the GCN class. Lemma G.2 yields that the predictor $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\cdot)$ generated at the last layer is an empirical risk minimizer in the function class $\mathcal{F}_{\text{cgcn}}$. As such, by the theory of Rademacher complexity (see, e.g., (Bartlett and Mendelson, 2002)), we obtain that:
@@
where $c$ is a universal constant a and we used the assumption that $J$ is $M$-Lipschitz. By Lemma G.1, we have that $\inf_{\Phi\in\mathcal{F}_{L,cgcn}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]\leq\inf_{\Phi\in\mathcal{F}_{L,gcn}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]$. Since $\mathcal{F}_{\text{cgcn}}=\mathcal{F}_{1,cgcn}$ due to $L=1$, this means that $\inf_{\Phi\in\mathcal{F}_{\text{cgcn}}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]\leq\inf_{\Phi\in\mathcal{F}_{\text{gcn}}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]$. Combined with (43), we obtain the desired.
Remark G.4.
The constant $C_{\sigma,L}(\mathcal{B}_{L})$ depends on the convergence rate of the polynomial expansion of the activation function $\sigma$ (See Appendix C). Algorithmically, knowing the activation function is not required to run Algorithm 2. Theoretically, however, the choice of $\sigma$ matters for Theorem 4.7 when comparing the predictor produced by Algorithm 2 against the best GCN, as it affects the representation power of such GCN. If a GCN with an activation function $\sigma$ performs well, CGCN will similarly generalize strongly.
Appendix H ADDITIONAL EXPERIMENTAL DETAILS
H.1 Dataset Descriptions
Table 2 provides a summary of the statistics and characteristics of datasets used in this paper. All datasets are accessed using the TUDataset class (Morris et al., 2020) in PyTorch Geometric (Fey and Lenssen, 2019) (MIT License).